# Retrieving Evidence from EHRs with LLMs: Possibilities and Challenges

### v2

@@ 3 Data
For evaluation we worked with radiologists (specializing in neuroimaging) from the Brigham and Women’s Hospital in Boston (BWH). For experiments, we used a private dataset from this hospital and the publicly available MIMIC-III (Johnson et al., 2016) dataset, to ensure that our findings are robust and (partially) reproducible.
BWH dataset comprises patients admitted to the Emergency Room (ER) of BWH between 2010 and 2015 along with clinical notes including: cardiology, endoscopy, operative, pathology, pulmonary, radiology reports, and discharge summaries. We sampled patients who underwent brain imaging within 48 hours of their ER visit because they are likely to have undetermined diagnoses. We are interested in scenarios where patients are associated with a large volume of EHR data, so we included patients with $\geq$10 EHR notes.
@@
Very Useful
intracranial hemorrhage
Recent fossa surgery and now on anticoagulants
Surgery in the brain inevitably leaves some hemorrhage. Anticoagulants increase the risk of hemorrhage. ‘Recent surgery’ and ‘anticoagulants’ make hemorrhage highly likely.
Useful
cerebral infarction
There is calcified thrombus obstructing the origins of the M2 branches
‘Thrombus’ is diagnostic of infarction, which is very useful information. But ‘calcified thrombus’ implies chronicity, so the thrombus could have been present for a long time and there may not be an acute infarction at this time.
Partially Useful
chemoradiation necrosis
The patient is at risk of chemoradiation necrosis due to her history of seizures and brain abscess, which may have caused damage to the brain tissue. Additionally, her use of concurrent Temodar and involved field radiation during her treatment may have further increased her risk.
History of seizures and brain abscess are not relevant to chemoradiation necrosis. Concurrent Temodar use and involved field radiation is useful information.
Weak Correlation
pneumocephalus
patient was involved in a motorcycle accident
A traumatic head injury is an important risk factor of pneumocephalus. A motorcycle accident increases the likelihood of a head injury.
Not Useful
small vessel disease (SVD)
patient is at risk of endocarditis
Not helpful in diagnosing SVD.
Hallucination
intracranial hemorrhage
patient has a brain tumor
Not present in the note.
@@
For expert evaluation, one of the collaborating radiologists identified all diagnoses discussed in the Findings and Impressions sections of the radiology reports of 10 patients from each dataset (excluding MIMIC-III patients from the pilot study). [^3] While this is a relatively small number of patients, we emphasize that manual evaluation is expensive: Radiologists on our team spent $\sim$9 hours manually assessing outputs. Then, for each diagnosis, we retrieved supporting evidence from all patient notes using the zero-shot prompting strategy from Section 2. Three collaborating radiologists then manually assessed each retrieved piece of evidence.
@@ 4.2 Weakly correlating evidence
CAD (s/p stents x 2, $>2$ MIs, on Coumadin INR=1.9) hx of $>3$ TIAs in past 2.5 yrs multiple AAAs (largest last measured at 5.5 cm, surg intervention held 2/2 cardiac status
IMMUNIZATIONS: INFLUENZA VACCINE (INACTIVATED) IM Given [DATE] ALLERGY: AMOXICILLIN ADMIT DIAGNOSIS: Stroke PRINCIPAL DISCHARGE DIAGNOSIS ;Responsible After Study for Causing Admission) same OTHER DIAGNOSIS;Conditions,Infections,Complications,affecting Treatment/Stay CAD (s/p stents x 2, $>2$ MIs, on Coumadin INR=1.9) hx of $>3$ TIAs in past 2.5 yrs multiple AAAs (largest last measured at 5.5 cm, surg intervention held 2/2 cardiac status OPERATIONS AND PROCEDURES: None.OTHER TREATMENTS/PROCEDURES (NOT IN O.R.)
The patient has a TBI
A/P- S/P REPAIR [**Doctor Last Name **] & LL ORTHOPEDIC INJURIES STABLE TBI W/CLOSE MONITORING FOR CHANGES STABLE LIVER LAC AT PRESENT SUCCESSFULL WEAN/EXTUBATION POST-OP PAIN CONT TO MONITOR PER ORDERS- Q2/HR NEURO & PERIPHERAL VASCULAR CHECKS…?
@@ 7 Discussion and Limitations
Our evaluation was limited in a few key ways. We enlisted radiologists to perform in-depth evaluation of a small number of instances, because evaluation is time consuming: We emphasize that this exercise required substantial allocation ($\sim$16 hours) of scarce expert time. We attempted to mitigate this via LLM-based automatic evaluation, performed at larger scale. However, our assessment of this strategy also relied on this relatively small annotated set and so may not generalize. Another limitation here is that we considered only two LLMs (specifically, FLAN-T5 and Mistral-Instruct): Other LLMs might, naturally, perform better or worse. Finally, we did not extensively iterate on the prompts used, and this too could substantially affect results.
@@ A.2 Few-shot prompting
pneumocephalus
head injury, skull fracture, neurosurgical procedures, sinus or mastoid surgery, meningitis, cerebrospinal fluid leak, barotrauma, diving or scuba diving accidents, iatrogenic causes, such as lumbar puncture or spinal anesthesia
stroke
hypertension, smoking, diabetes, obesity, sedentary lifestyle, high cholesterol levels, atrial fibrillation, family history of stroke, previous history of stroke, excessive alcohol consumption, drug abuse.
intracranial hemorrhage
hypertension, aneurysms, arteriovenous malformations, blood clotting disorders, trauma, drug abuse, liver disease, brain tumor, stroke, coagulopathy
brain tumor
progression genetics, exposure to ionizing radiation, family history of brain tumors, certain hereditary conditions, weakened immune system, previous history of brain tumor.
intracranial hypotension
obesity, connective tissue disorders, previous spinal or cranial surgery, leaking cerebrospinal fluid, spinal epidural anesthesia, lumbar puncture or spinal tap
@@
intracranial hemorrhage
patient had multiple cardiac surgeries
Multiple cardiac surgeries may suggest anticoagulation or underlying cardiac dysfunction which could in turn predispose the patient to intracranial hemorrhage.
intracranial hypotension
The patient has a ventriculoperitoneal shunt.
A ventriculoperitoneal shunt (VPS) is a surgical device used to relieve intracranial pressure by draining excessive cerebrospinal fluid. Having a VPS catheter may increase the risk of intracranial hypotension due to over drainage.
craniopharyngioma
s/p resection X2, s/p VPS and panhypopitutiarism with second resection
Panhypopituitarism and the fact that something was removed through surgery suggests there was a tumor involving the sella which may or may not have been craniopharyngioma.

### v3

@@ Abstract
Data and Code Availability
We describe the data used for evaluation in §3. Briefly, we evaluate our approach using two datasets: (1) MIMIC-III dataset, (Johnson et al., 2016b), which is available on PhysioNet (Johnson et al., 2016a); and (2) EHR notes of patients admitted to the Emergency Room of Brigham and Women’s Hospital (BWH) in Boston, MA, USA, between 2010 and 2015. Our code and data are available at https://github.com/hibaahsan/chil_diagnosis_evidence/.
Institutional Review Board (IRB)
This retrospective medical records research was approved by the Mass General Brigham (MGB) IRB with a waiver of requirement for informed consent.
@@ 3 Data
For evaluation, we collaborated with radiologists (specializing in neuroimaging) from the Brigham and Women’s Hospital in Boston (BWH). Three radiologists with 25,15, and 8 years of experience, respectively, participated in the evaluation. One of them (25 years of experience) had prior experience with LLM projects while the other two did not. For experiments, we used a private dataset from this hospital and the publicly available MIMIC-III (Johnson et al., 2016b) dataset, to ensure that our findings are robust and (partially) reproducible.
BWH dataset comprises patients admitted to the Emergency Room (ER) of BWH between 2010 and 2015 along with clinical notes including: cardiology, endoscopy, operative, pathology, pulmonary, radiology reports, and discharge summaries. We sampled patients who underwent brain imaging within 48 hours of their ER visit. It is typically in the ER that evidence can be beneficial as the initial diagnosis is undetermined in most ER cases. Clinicians attend to several patients in one shift and have to go through often previously unknown patient history to come up with a diagnosis (Murray et al., 2021). Without this constraint, the probability of the diagnosis already being determined in the past would be higher and would trivialize the problem. We are interested in scenarios where patients are associated with a large volume of EHR data, so we included patients with $\geq$10 EHR notes.
@@
Very Useful intracranial hemorrhage Recent fossa surgery and now on anticoagulants Surgery in the brain inevitably leaves some hemorrhage. Anticoagulants increase the risk of hemorrhage. ‘Recent surgery’ and ‘anticoagulants’ make hemorrhage highly likely.
Useful cerebral infarction There is calcified thrombus obstructing the origins of the M2 branches ‘Thrombus’ is diagnostic of infarction, which is very useful information. But ‘calcified thrombus’ implies chronicity, so the thrombus could have been present for a long time and there may not be an acute infarction at this time.
Partially Useful chemoradiation necrosis The patient is at risk of chemoradiation necrosis due to her history of seizures and brain abscess, which may have caused damage to the brain tissue. Additionally, her use of concurrent Temodar and involved field radiation during her treatment may have further increased her risk. History of seizures and brain abscess are not relevant to chemoradiation necrosis. Concurrent Temodar use and involved field radiation is useful information.
Weak Correlation pneumocephalus patient was involved in a motorcycle accident A traumatic head injury is an important risk factor of pneumocephalus. A motorcycle accident increases the likelihood of a head injury.
Not Useful small vessel disease (SVD) patient is at risk of endocarditis Not helpful in diagnosing SVD.
Hallucination intracranial hemorrhage patient has a brain tumor Not present in the note.
@@
For expert evaluation, one of the collaborating radiologists identified all diagnoses discussed in the Findings and Impressions sections of the radiology reports of 10 patients from each dataset (excluding MIMIC-III patients from the pilot study). [^3] While this is a relatively small number of patients, we emphasize that manual evaluation is expensive: Radiologists on our team spent $\sim$9 hours manually assessing outputs. Then, for each diagnosis, we retrieved supporting evidence from all patient notes using the zero-shot prompting strategy from Section 2. The three collaborating radiologists then manually assessed each retrieved piece of evidence.
@@ 4.2 Weakly correlating evidence
CAD (s/p stents x 2, $>2$ MIs, on Coumadin INR=1.9) hx of $>3$ TIAs in past 2.5 yrs multiple AAAs (largest last measured at 5.5 cm, surg intervention held 2/2 cardiac status IMMUNIZATIONS: INFLUENZA VACCINE (INACTIVATED) IM Given [DATE] ALLERGY: AMOXICILLIN ADMIT DIAGNOSIS: Stroke PRINCIPAL DISCHARGE DIAGNOSIS ;Responsible After Study for Causing Admission) same OTHER DIAGNOSIS;Conditions,Infections,Complications,affecting Treatment/Stay CAD (s/p stents x 2, $>2$ MIs, on Coumadin INR=1.9) hx of $>3$ TIAs in past 2.5 yrs multiple AAAs (largest last measured at 5.5 cm, surg intervention held 2/2 cardiac status OPERATIONS AND PROCEDURES: None.OTHER TREATMENTS/PROCEDURES (NOT IN O.R.)
The patient has a TBI A/P- S/P REPAIR [**Doctor Last Name **] & LL ORTHOPEDIC INJURIES STABLE TBI W/CLOSE MONITORING FOR CHANGES STABLE LIVER LAC AT PRESENT SUCCESSFULL WEAN/EXTUBATION POST-OP PAIN CONT TO MONITOR PER ORDERS- Q2/HR NEURO & PERIPHERAL VASCULAR CHECKS…?
@@ 7 Discussion and Limitations
Our evaluation was limited in a few key ways. We enlisted radiologists to perform in-depth evaluation of a small number of instances, because evaluation is time consuming: We emphasize that this exercise required substantial allocation ($\sim$16 hours) of scarce expert time. We attempted to mitigate this via LLM-based automatic evaluation, performed at larger scale. However, our assessment of this strategy also relied on this relatively small annotated set and so may not generalize. Another limitation here is that we considered only two LLMs (specifically, FLAN-T5 and Mistral-Instruct): Other LLMs might, naturally, perform better or worse. In addition, we did not investigate the fairness implications of our work. However, the small size of our expert-annotated sample and the inherently small samples of underrepresented groups limits our ability to meaningfully assess this. We leave the detailed analysis needed to determine if there are significant differences to future work. Finally, we did not extensively iterate on the prompts used, and this too could substantially affect results.
8 Acknowledgment
We acknowledge partial funding for this work by National Library of Medicine of the National Institutes of Health (NIH) under award numbers R01LM013772 and R01LM013891. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@ A.2 Few-shot prompting
pneumocephalus head injury, skull fracture, neurosurgical procedures, sinus or mastoid surgery, meningitis, cerebrospinal fluid leak, barotrauma, diving or scuba diving accidents, iatrogenic causes, such as lumbar puncture or spinal anesthesia
stroke hypertension, smoking, diabetes, obesity, sedentary lifestyle, high cholesterol levels, atrial fibrillation, family history of stroke, previous history of stroke, excessive alcohol consumption, drug abuse.
intracranial hemorrhage hypertension, aneurysms, arteriovenous malformations, blood clotting disorders, trauma, drug abuse, liver disease, brain tumor, stroke, coagulopathy
brain tumor progression genetics, exposure to ionizing radiation, family history of brain tumors, certain hereditary conditions, weakened immune system, previous history of brain tumor.
intracranial hypotension obesity, connective tissue disorders, previous spinal or cranial surgery, leaking cerebrospinal fluid, spinal epidural anesthesia, lumbar puncture or spinal tap
@@
intracranial hemorrhage patient had multiple cardiac surgeries Multiple cardiac surgeries may suggest anticoagulation or underlying cardiac dysfunction which could in turn predispose the patient to intracranial hemorrhage.
intracranial hypotension The patient has a ventriculoperitoneal shunt. A ventriculoperitoneal shunt (VPS) is a surgical device used to relieve intracranial pressure by draining excessive cerebrospinal fluid. Having a VPS catheter may increase the risk of intracranial hypotension due to over drainage.
craniopharyngioma s/p resection X2, s/p VPS and panhypopitutiarism with second resection Panhypopituitarism and the fact that something was removed through surgery suggests there was a tumor involving the sella which may or may not have been craniopharyngioma.